# Is This Collection Worth My LLM’s Time? Automatically Measuring Information Potential in Text Corpora

### v1

@@ A.2.2 Information Potential
(a) Llama 70B Baseline
@@
(b) Llama 70B Wikipedia
IP: 0.136
(c) Llama 70B EPFL
IP: 0.229
(d) GPT-4o Wikipedia
IP: 0.110
(e) GPT-4o EPFL
IP: 0.211
(f) Legend
@@
(a) Baseline
(b) Wikipedia
(c) EPFL
(d) Legend
@@
Information Potential: The value computed with Equation 3
@@ A.2.3 Qualitative Analysis of EPFL MCQ Dataset
[…] Symmetric joins require that all relations be materialized and hence incur materialization overhead. To reduce the overhead, RouLette materializes only tuples that can form output tuples for their query-set. We call this symmetric join pruning […] Novel Findings
@@
[…] its clinical outcome was rather disappointing 4 . One of reasons for this seems to be the limiting number of available endogenous mesenchymal progenitor cells (MPCs) that can give rise to bone cells. […]. Hence, there is a clear clinical need for implants that augment the homing/recruitment of endogenous MPCs to fracture sites […] Novel Findings
@@
[…] The higher average THC concentrations levels with time was caused by the gradual accumulation of carbonates on ceria sites during the periodic oscillations. […] Novel Findings

### v2

@@ A.2.2 Information Potential
(a) Llama 70B EPFL
IP: 0.229
(b) GPT-4o EPFL
IP: 0.211
(c) GPT-4o Venice
IP: 0.265
(d) Llama 70B Wikipedia-EPFL
IP: 0.136
(e) GPT-4o Wikipedia-EPFL
IP: 0.110
(f) GPT-4o Wikipedia-Venice
IP: 0.180
(g) Llama 70B Baseline
@@
(h) Legend
@@
001010101020202020303030304040404050505050606060600.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Cutoff Percentile ValuesFraction of Correct ResponsesBaselineWikipedia-EPFLEPFLWC 4xNC 4xIP
@@
IP Score: The Information Potential computed with Equation 3
@@ A.2.3 Qualitative Analysis of EPFL MCQ Dataset
[…] Symmetric joins require that all relations be materialized and hence incur materialization overhead. To reduce the overhead, RouLette materializes only tuples that can form output tuples for their query-set. We call this symmetric join pruning […] Unique Mentions
@@
[…] its clinical outcome was rather disappointing 4 . One of reasons for this seems to be the limiting number of available endogenous mesenchymal progenitor cells (MPCs) that can give rise to bone cells. […]. Hence, there is a clear clinical need for implants that augment the homing/recruitment of endogenous MPCs to fracture sites […] Unique Mentions
@@
[…] The higher average THC concentrations levels with time was caused by the gradual accumulation of carbonates on ceria sites during the periodic oscillations. […] Unique Mentions